The odds against Thunderbolt winning the Birmingham Derby are 11:3. Find the probability that Thunderbolt wins.
A) 3/11 B) 3/14 C) 8/11 D) 8/14 E) 3/8

Answer by stanbon(75887)

The odds against Thunderbolt winning the Birmingham Derby are 11:3. Find the probability that Thunderbolt wins.
---
Odds against event A = [P(A')]/[P(A)]
---
Equation:
[P(A')]/[P(A)]= 11/3 = [11/14]/[3/14]
P(A) = 3/14
======================
Cheers,
Stan H.

A) 3/11 B) 3/14 C) 8/11 D) 8/14 E) 3/8